OpenStudy (anonymous):

How exactly would I do this? 4t-3st=7+s Any help would be appreciated

OpenStudy (austinl):

What are we doing? Solving for a particular variable? Simplifying?

OpenStudy (anonymous):

Solve for s

OpenStudy (phi):

you start with 4t-3st=7+s you want to get all the terms with s on the same side add 3st to both sides 4t -3st +3st = 7 + s + 3st simplify that to 4t = 7 + s + 3st "move" the 7 to the left side, by adding -7 to both sides 4t - 7 = s + 3st

OpenStudy (phi):

the next step is factor s from each term on the right side can you do that ??

OpenStudy (austinl):

\(-3st+4t=s+7\) Isolate s to the left hand side. \(s(-1-3t)=7-4t\) \(s=\dfrac{7-4t}{-1-3t}\) *DISCLAIMER* I believe this is how you do it, I am quite rusty at this.